### v2

# Word Definitions from Large Language Models

Abstract
Dictionary definitions are historically the arbitrator of what words mean, but this primacy has come under threat by recent progress in NLP, including word embeddings and generative models like ChatGPT. We present an exploratory study of the degree of alignment between word definitions from classical dictionaries and these newer computational artifacts. Specifically, we compare definitions from three published dictionaries to those generated from variants of ChatGPT. We show that (i) definitions from different traditional dictionaries exhibit more surface form similarity than do model-generated definitions, (ii) that the ChatGPT definitions are highly accurate, comparable to traditional dictionaries, and (iii) ChatGPT-based embedding definitions retain their accuracy even on low frequency words, much better than GloVE and FastText word embeddings.
1 Introduction
Many generations of readers have relied on the notion that the meaning of a word is what it says in the dictionary. Definitions of the primary and alternate senses of words have appeared in printed books well before Samuel Johnson’s famous A Dictionary of the English Language in 1755. Early western lexicography evolved from glosses and glossaries. American English was later standardized by Noah Webster in American Dictionary of the English Language in 1828.
But the primacy of dictionary definitions as the arbitrator of word meanings has come under threat by recent progress in natural language processing. Vector representations (word embeddings) have proven more valuable computationally than handcrafted semantics or definitions in describing the effective meaning of vocabulary words in a given language. And generative dialog systems like ChatGPT will happily produce full text definitions of every word when asked; indeed multiple versions of these definitions in response to different prompts.
This paper is an exploratory study of the degree of alignment between word definitions from classical dictionaries and these newer computational artifacts. There are several distinct questions we address here:
  • How consistent is the representation of semantics from classical dictionary definitions with those of word embeddings and generative language models?
  • How quickly are the definitions generated by LLMs improving as the technology advances?
  • LLMs have been described as “stochastic parrots” Bender et al. (2021). To what extent are the definitions they produce mere repetitions or even plagiarisms of classical dictionary definitions?
Our methodology is as follows. For each of over 2,500 carefully selected words, we extracted their handcrafted definitions in each of three different dictionaries (WordNet, Merriam-Webster, and Random House/Dictionary.com), and also vector representations from two prominent collections of word embeddings (FastText and Glove). We also asked two versions of ChatGPT (3.5 and 4.0) for the definitions of every word, each with two different forms of prompts. The first prompt is "what is the meaning of this word?", while the second prompt is "define this word." We measure the distances between these representations using several techniques, including vector distance/similarity, edit distance, and neighborhood correlations.
Our primary observations include:
  • The surface form of generated definitions differ substantially from those of published dictionaries – Although ChatGPT presumably trained in part on the handcrafted definitions in our study, little-to-no unexpected trace of these texts remain in the generated text. Indeed, we find roughly twice as many long exact matches between traditionally published dictionary definitions than we observe between published and generated texts.
  • ChatGPT generated word definitions are highly accurate, consistent with published dictionaries – We evaluate SBERT definitions embeddings to identify candidates for mismatching definitions across sources. Of the 50 most distant pairs between Merriam-Webster and GPT4 definitions, human evaluation confirms that GPT4 recognized 48 compared to 34 for GPT3, and that all but one of these definitions matched the primary sense of the published dictionary.
  • The consistency of LLM-generated definitions is relatively independent of frequency, unlike word embeddings – We propose a new average distance correlation metric to compare word embeddings in different dimensional spaces, and use it to study both traditional word embeddings (GloVE and FastText) and sentence embeddings of both dictionary and generated definitions. Our results suggest that SBERT embeddings of definitions (synthesized or published) may create more accurate word embeddings for low frequency words than traditional methods.
This paper is organized as follows. Section 2 presents related work on dictionaries, definitions, and semantic representation. The dictionaries and generative models we use to formulate our dataset are described in Section 3. The question of just how originial generated definitions are is considered in Section 4. We use vector representations to quantify the similarity of definitions by source in Section 5, and between definitions and word embeddings in Section 6. We conclude with directions for future work in Section 7.
2 Related Work
Large language models (LLM) recently burst into popular consciousness through the phenomenon of ChatGPT, which reached 100 million users in only three months Milmo (2023), offering the prospect of knowledge systems that can write professional-quality documents. Generative Pre-trained Transformer (GPT) models are a series of large language models (LLMs) developed by OpenAI. Each model in the series Radford and Narasimhan (2018) Radford et al. (2019) Brown et al. (2020) OpenAI (2023) is trained on a larger corpus of text and achieves better results on natural language processing (NLP) tasks compared to its predecessors.
ChatGPT promises to have a major impact on many fields, including healthcare Scerri and Morin and education Han et al. (2023); Mogali . The models have also been used to assess translation quality Kocmi and Federmann (2023) and the nature of human personalities Rao et al. (2023).
2.1 LLMs and Prompt Engineering
One aspect of our work is assessing how different prompts to LLMs (here, for word definitions) yield responses of different quality. Prompt engineering has been studied for Human-Computer Interaction Oppenlaender (2022), NLP Liu et al. (2021); Wang et al. (2023), and tool learning based on foundation models Qin et al. (2023). Prompt design helps overcome model bias Rao et al. (2023). Chen et al. identify the best prompt to induce correct answer from pre-trained language models at a low cost Chen et al. (2023).
Chain-of-thought prompting Wei et al. (2023) decomposes a task into a sequence of steps. It has been used to allow LLMs to use tools such as a calculator and a chemical reaction predictor Inaba et al. (2023), to improve reasoning over long input documents Sun et al. (2023), to perform better on web navigation tasks Sridhar et al. (2023), and to achieve more correct answers through dynamic sampling Aggarwal et al. (2023).
2.2 Dictionaries and Definitions
Lexicographical works can be dated back to ancient Mesopotamia Veldhuis (2019) and ancient China Creamer (1992). The ancient Greeks also use glossaries to understand religious and literature works Lee (2022); Dolezal and Velde (2021). Medieval lexicographers use derivation and alphabetization techniques Weijers (1989), which lay the foundations for early modern dictionaries Considine (2008).
The English language was standardized through dictionaries by Samuel Johnson and Noah Webster Sokolowski (2020). The Oxford English Dictionary represents another milestone, with multiple editions created in the 19th and 20th centuries. Most dictionaries in the digital age now appear in electronic form Zimmer (2014).
Researchers have shown that typical native speakers of American English know an average of 42,000 words and 4,200 multiword expressions. This number ranges from 27,000 words for the lowest 5% to 52,000 words for the highest 5% Brysbaert et al. (2016).
2.3 Computational Use of Dictionaries
A lexicon with information about word meanings and usage is important to solve many NLP tasks. Doman-specific dictionaries for NLP have been developed over the years, initially through handcrafting and then automatic construction Riloff (1996); Lovis et al. (1998); Dolan et al. (1999); Przepiórkowski et al. (2014).
One interesting NLP problem tackled using dictionaries is word sense disambiguation (WSD), which explores the use of word sense in a particular context. Lexicographers study word senses in dictionaries Kilgarriff (1992) and suggest that computer readable dictionary corpora provide new opportunities for lexicography. Markowitz et al. Markowitz et al. (1986) and Montemagni et al. Montemagni and Vanderwende (1992) extract lexical information from implicit information in dictionary definitions. Dolan (1994) suggests that dictionary definitions are too fine-grained to use for WSD, and proposes a method to cluster word senses based on WordNet definitions and semantic relations. Pentheroudakis et al. (1994) identify morphologically related words in machine readable dictionaries (MRDs) and provide a certainty score for the extracted relations. Automatically gathered information from dictionaries have also been used to enrich thesauruses Oliveira and Gomes (2013).
2.4 Definitions and Word Embeddings
Word embeddings are a multi-dimensional vector representation of a word that conveys semantic and syntactic information. Words with similar meanings have embeddings closer to each other in the vector space. Popular word embeddings like Word2vec Mikolov et al. (2013), GloVe Pennington et al. (2014), and FastText Bojanowski et al. (2017) have been used for many NLP tasks.
Definition modeling seeks to generate word definitions given a word and its embedding Noraset et al. (2016). Definition generation solves the WSD problem by creating definitions instead of choosing a correct one Bevilacqua et al. (2021). Chang et al. analyze what contexualized word representations capture, and show these representations can be interpreted in a human-readable way Chang and Chen (2019). Huang et al. add a re-ranking mechanism to the decoder-encoder model to control specificity Huang et al. (2021). Liu et al use capsule networks first proposed in Sabour et al. (2017) to decompose word embeddings into morpheme-like vectors and reconstruct them into sense embeddings. Word embeddings have also been extracted from Urban dictionary Wilson et al. (2020a) and Twitter Wilson et al. (2020b).
3 Dataset
Rate limitations on using the LLM APIs prevented us from studying the definitions of all English-language words, so we were selective in what words we study. Starting from a lexicon containing 246,591 English words sorted by their frequency, we sample 3000 words representing 500 words in each of six frequencies tiers:
  • Common: rank 1 to 500, and 501 to 1000.
  • Less Common: 5000-5500, and 10000-10500.
  • Rare: 20000-20500, and 50000-50500.
We cleaned the dataset by removing any word whose definition is missing in any source. This leaves 2512 words with complete definitions remaining in our final dataset. Table 1 presents a breakdown of these words by frequency tier and part of speech (POS) tag, demonstrating that our sampled words are well distributed by both features. Nouns are the most common type of words in the dataset.
Common Less common Rare Total
NOUN 749 475 182 1406
VERB 179 149 53 381
ADJ 153 77 19 249
ADV 83 32 16 131
Other 55 0 290 345
Total 1219 733 560 2512
Table 1: Word frequency and POS tag summary.
For each selected lexicon word, we collect answers from both GPT3.5 and GPT4 using two different prompts, identified throughout the text as type 1 or 2 prompts:
  • Type 1: “What is the meaning of this word?"
  • Type 2: “Define this word."
This experimental design enables us to make a variety of comparisons between definitions produced (a) between different prompts on the same model, (b) two different models, namely GPT3.5 and GPT4.0, (c) comparisons between models and published dictionaries, (d) between common, moderate, and rare words, and (e) between different parts of speech.
To provide fair grounds for comparison, we extracted published definitions for each of these words from three online dictionaries, namely:
  • WordNet Miller (1995) (http://wordnetweb.princeton.edu/perl/webwn): This manually constructed database of word definitions and semantic relationships among words. It is highly accurate and serves as a useful tool in the field of NLP and linguistics.
  • Merriam-Webster dictionary Merriam-Webster Online (2023) (https://www.merriam-webster.com): The Merriam-Webster.com site is a commercial digital dictionary which is regularly updated.
  • Dictionary.com (https://www.dictionary.com): An online dictionary based on the Random House Unabridged Dictionary, and is updated regularly by editors.
The length distribution of these sources is summarized in Table 2, including mean and min/max lengths and standard deviation. Observe that GPT4 definitions are generally a little longer than those produced by GPT3.5, but within a given model the two prompts produce similar length texts. The correlation of the lengths of the two GPT3 definitions is 0.61, and that of the two GPT4 definitions is 0.74.
Two of the published dictionaries (Merriam-Webster and Dictionary.com) produce substantially longer definitions than WordNet and GPT3.5.
Source
N
Avg
Length
Min
Max
$\sigma$
GPT3-1
2459 35.97 3 253 20.21
GPT3-2
2488 43.09 3 248 28.65
GPT4-1
2488 132.31 20 386 64.97
GPT4-2
2489 133.77 28 466 65.26
Wordnet
2247 37.25 1 416 45.36
Merriam-W.
2231 132.43 2 1957 179.59
Dict.com
2398 223.11 3 3049 328.89
Table 2: Length summary for collected dictionary datasets, showing the mean, min, max and standard deviation for each source, plus the number of queried lexicon terms defined within each sources (out of 3000).
4 Does ChatGPT Plagiarize?
Although neural network-based generative models synthesize response texts instead of explicitly cutting-and-pasting from training texts, this does not mean they cannot plagiarize, say through overfitting model parameters. Dictionary definitions provide an interesting domain to assess this possibility: they are precisely written texts that are readily available in training data. There is a long history of dictionaries plagiarizing each other Cooper (1962), and commonly include definitions of fake words (such as Mountweazel or jungftak) in order to catch wholesale pirating by other works Williams (2016).
Various plagiarism detection techniques Foltýnek et al. (2020); Zhou et al. (2022) can be used to check if GPT models are plagiarizing. The originality of essays produced by ChatGPT can be evaluated using plagiarism detection tools Khalil and Er (2023). The result indicates that ChatGPT output is complex enough to avoid being caught, and that ChatGPT itself performs better on determining whether the text is generated by the model or not.
We investigate this issue by identifying the longest common substring between the ChatGPT-generated definition for a given word and the corresponding definition in a published dictionary. Table 3 reports the words which contain the longest case-insensitive match between each model/dictionary pair. These examples are quite compelling, with common sequences as long as 17 words capturing the heart of the definition.
Source GPT Word Longest Common Substring
Wordnet 3-1 clinic a medical establishment run by a group of medical
3-2 minutes a unit of time equal to 60 seconds or 1/60th of an hour
4-1 forerunner that precedes and indicates the approach of something or someone
4-2 nodule ally harder than the surrounding rock or sediment
Merriam -Webster 3-1 kelvin the base unit of temperature in the International System of Units
3-2 town ally larger than a village but smaller than a city
4-1 atheism a philosophical or religious position characterized by disbelief in the existence of a god or any gods
4-2 econometric tical methods to the study of economic data and problems
Dictionary .com 3-1 letter addressed to a person or organization and usually transmitted by mail
3-2 delta the fourth letter of the Greek alphabet ($\Delta$, $\delta$)
4-1 back he rear part of the human body, extending from the neck to the lower end of the spine
4-2 compared o examine (two or more objects, ideas, people, etc.) in order to note similarities and differences
Table 3: The longest common substring between GPT-model generated definitions and published dictionaries. These matching phrases often capture the primary sense of the underlying word.
These examples are provocative, but not necessarily damning. Determining a meaningful expected longest common substring length in our instance is challenging, for several reasons. For randomly generated sequences, statisticians have proven this quantity grows logarithmically in the length of the sequences Arratia and Waterman (1985); Barros et al. (2019). But natural language text is far from random, and uses a large vocabulary as opposed to a constant-sized alphabet. Further, text with a high semantic similarity (two definitions of the same word) should share greater surface similarity than mismatched definitions. The length of the definitions matter, in a non-trivial way: longer, more detailed definitions should be expected to contain longer matches than briefer descriptions.
To assess whether these common phrases reflect untoward plagiarism or instead naturally constrained word choices in precise definitions, we compare the degree of borrowing between published dictionaries with what one sees generated by models. Standards for acceptable borrowing in published dictionaries are established by social convention and copyright law. Do LLMs borrow more from published dictionaries than dictionaries borrow from each other?
(a) Match length distributions for dictionaries and GPT models.
(b) Average match length distributions against dictionaries.
Figure 1: Frequency distribution of number of matched words, among dictionary pairs and GPT models (top), and the average number of matched words among dictionaries and between GPT model and dictionaries (bottom). GPT models exhibit substantially more borrowing than published dictionaries.
Figure 1 presents the frequency distributions of match length between all three pairs of published dictionaries, and the two prompts for each model. They show that each pair of dictionaries share a small tail of long common substrings in their definitions, with substantially greater sharing between the two definitions from GPT-3 and GPT-4. There is nothing untoward about this, just that the generative models repeat language in the two equivalent definitions. Surprisingly, GPT-4 repeats itself substantially more frequently than GPT-3, perhaps an artifact from efforts to reduce model hallucination.
The average length of common strings, and frequencies of long matches reported in Table 4 shows more aggressive borrowing between Merriam-Webster and Dictionary.com than any other pairing of published dictionaries or dictionary-model pairs. In particular, the 155 long matches (of length $\geq 5$) observed between Merriam-Webster and Dictionary.com were roughly twice as many as between either of these dictionaries and any model. There seems no real evidence that GPT models unfairly replicate published training data, despite the long matches reported in Table 3.
Matched Pair
Average
longest
match
Count
matches
len5absent5$\geq$ 5≥ 5
Wordnet / Merriam-Webster 1.32 98
Wordnet / Dictionary.com 1.13 77
Merriam-Webster / Dictionary.com 1.54 155
GPT3-1 / Wordnet 1.55 99
GPT3-2 / Wordnet 1.59 97
GPT4-1 / Wordnet 1.65 93
GPT4-2 / Wordnet 1.62 91
GPT3-1 / Merriam-Webster 1.23 80
GPT3-2 / Merriam-Webster 1.25 75
GPT4-1 / Merriam-Webster 1.20 77
GPT4-2 / Merriam-Webster 1.18 70
GPT3-1 / Dictionary.com 1.00 69
GPT3-2 / Dictionary.com 1.00 71
GPT4-1 / Dictionary.com 1.04 65
GPT4-2 / Dictionary.com 1.04 75
GPT3-1 / GPT3-2 3.09 592
GPT4-1 / GPT4-2 4.00 711
Table 4: Average length and number of long matches ($\geq 5$ words) between definition sources.
5 Are GPT Definitions Consistent with Dictionaries?
Model-generated word definitions aim to capture the same essential meaning as published dictionaries, while avoiding textual equivalence. In order to obtain quantitative information about the relative quality of published and generated definitions, we consider three distinct distance or similarity functions between short texts—here definitions of the same word from different sources. Two of these are based on the distance between embeddings, while the third works in the full text space of the definitions:
  • BERT Reimers and Gurevych (2019) – a modification of BERT that uses Siamese network structures to extract sentence-level representation using triplet loss.
  • MPnet Song et al. (2020) – a pre-trained model that combines permuted language modeling and the use of auxiliary position information to derive feature embedding.
  • Edit Distance – a string metric that represents the minimum number of insertions, deletions, or replacements required to transform one string into another. Here the edit distance is normalized by dividing itself by the maximum length of the two strings.
We calculate the distance between definitions provided by GPT models and online dictionaries. Table 5 shows the cosine distance between each pair of online dictionaries based on sentence embeddings and edit distance. The definitions from ChatGPT appear closer to Wordnet and Merriam-Webster, while definitions from GPT4 are closer to those on Dictionary.com.
Dictionary GPT model BERT MPNet
Edit
Dist.
Wordnet GPT3-1 0.33 0.43 4.72
GPT3-2 0.33 0.44 4.81
GPT4-1 0.37 0.49 5.01
GPT4-2 0.35 0.45 5.08
Merriam -Webster GPT3-1 0.31 0.38 4.49
GPT3-2 0.30 0.39 4.50
GPT4-1 0.33 0.43 4.53
GPT4-2 0.31 0.40 4.59
Dictionary .com GPT3-1 0.36 0.31 4.92
GPT3-2 0.35 0.31 4.86
GPT4-1 0.33 0.32 4.59
GPT4-2 0.31 0.28 4.63
Table 5: Distance between online dictionaries and definitions provided by GPT models (cosine).
Similarly, Table 6 reports the distance between each pair of GPT model definitions. For both GPT-3.5 and GPT-4, the definitions produces by the same model with different prompts are closest by all three distance metrics. Both BERT and MPNet-based distances perform well in capturing the similarity of definition pairs. We use cosine distance of BERT embeddings as our distance function in the rest of this paper, although similar results follow from MPNet.
Model 1 Model 2 BERT MPNet
Edit
Dist.
GPT3-1 GPT3-2 0.18 0.18 3.96
GPT4-1 0.28 0.24 4.79
GPT4-2 0.25 0.19 4.86
GPT3-2 GPT3-1 0.18 0.18 3.96
GPT4-1 0.26 0.22 4.61
GPT4-2 0.22 0.17 4.65
GPT4-1 GPT3-1 0.28 0.24 4.79
GPT3-2 0.26 0.22 4.61
GPT4-2 0.17 0.12 3.84
GPT4-2 GPT3-1 0.25 0.19 4.86
GPT3-2 0.22 0.17 4.65
GPT4-1 0.17 0.12 3.84
Table 6: Distance between definitions generated from GPT models (cosine). The two distinct definitions produced by the same version of GPT are closer than those generated by different models.
5.1 POS Analysis
A natural hypothesis is that certain classes of words are easier to generate reliable definitions for than others. Nouns represent objects, which might appear to be easier to precisely than descriptive words like adjectives and adverbs. We obtain a dominant POS tag on each word using the NLKT library Bird et al. (2009). Table 7 presents the average cosine distance error between generated and published definitions, for each model and POS type. Curiously, GPT3.5 proves most accurate on descriptive words, which GPT4 produces its best definitions for nouns. The results are consistent across all three published dictionaries, and indistinguishable for Type 1 and Type 2 prompts.
Model POS WordNet
Merriam-
Webster
Dictionary.
com
GPT3-1 NOUN 0.32 0.30 0.35
VERB 0.32 0.34 0.42
ADJ 0.33 0.28 0.35
ADV 0.29 0.28 0.41
GPT3-2 NOUN 0.33 0.30 0.34
VERB 0.32 0.33 0.40
ADJ 0.36 0.27 0.34
ADV 0.29 0.27 0.39
GPT4-1 NOUN 0.36 0.32 0.31
VERB 0.37 0.35 0.35
ADJ 0.40 0.33 0.33
ADV 0.45 0.40 0.37
GPT4-2 NOUN 0.35 0.30 0.30
VERB 0.35 0.33 0.34
ADJ 0.38 0.31 0.31
ADV 0.41 0.35 0.35
Table 7: Cosine distance based on BERT for different POS tags.
5.2 Word Frequency Analysis
There are natural but contradictory hypotheses to govern how the quality of LLM-generated definitions should vary as a function of the relative frequency of each word. One may speculate that the most common words are hardest to define, because of functional forms like prepositions, and that they are more likely to be enriched with multiple senses. But it is equally reasonable to think that low frequency words will be most difficult for generative models to understand, as they are seen least frequently in training data.
To resolve this debate, we compare the average cosine distance between BERT-encoded dictionary and generated definitions, partitioned by class into words of high frequency, moderate frequency, and rare words. The results in Table 8 show words of middle frequency produce the best dictionary-generated for all models, across all dictionaries. The differences between frequency tiers is generally quite modest, generally less than the impact of model version. GPT-3.5 appears to generate slightly better definitions than the later GPT-4 by this metric, while the choice of prompt has little effect on the accurracy of resdult from either model.
Model Frequency
Word
Net
Merriam-
Webster
Dictionary.
com
GPT3-1 high 0.35 0.34 0.42
medium 0.28 0.28 0.31
low 0.32 0.28 0.30
GPT3-2 high 0.35 0.33 0.41
medium 0.29 0.27 0.30
low 0.34 0.29 0.30
GPT4-1 high 0.40 0.37 0.37
medium 0.34 0.29 0.28
low 0.37 0.29 0.28
GPT4-2 high 0.37 0.34 0.35
medium 0.33 0.28 0.28
low 0.39 0.30 0.28
Table 8: Cosine distance based on BERT for different word frequency.
5.3 Are GPT-Generated Definitions Accurate?
Researchers have found that ChatGPT can easily generate answers with complete assurance, even sometimes the answer is wrong Lehnert (2023). To evaluate the correctness of GPT-generated definitions, we manually compare the definitions of the 50 words with the largest Euclidean distance between embedding definitions between Merriam-Webster and GPT-4.
We tabulate the words with the largest such distances in Table 9. Of the fifty words we manually evaluate, the biggest distances occur when models admit they do not know the definitions of generally obscure words (often proper names and abbreviations) that appeared in Merriam-Webster. Only two of the fifty words (cordier and imon) were unknown to GPT-4, compared to 16 words GPT-3 did not understand. Of the 48 words for which GPT-4 ventured a definition, we deemed all of them similar except for “Acton”, which denoted a place in GPT-4 and a person in Merriam-Webster. We conclude that the model generated definitions are generally of high quality, consistent with those from published dictionaries.
Word
Definition from GPT3
Definition from GPT4
Definition in Merriam-Webster
paget I’m sorry, as an AI language model, I cannot define the word "paget" accurately without any context. "Paget" is primarily known as a surname of Norman origin and can be seen in various contexts. Sir James 1814–1899 English surgeon and pathologist.
acton I’m sorry, but "acton" is not a word in the English language. "Acton" is primarily known as a place name, particularly in England. It’s a district in the West London, in the Borough of Ealing. 1st Baron 1834–1902 John Emerich Edward Dalberg-Acton, English historian.
nov I’m sorry, I cannot define the word "nov" as it does not exist or have any clear meaning in the English language. "Nov" is most commonly recognized as an abbreviation for the month of November. November.
cordier As an AI language model, I cannot determine the context of the word "cordier". I’m sorry, but as of my knowledge cutoff in September 2021, "cordier" doesn’t have a specific definition in English. It seems like it could be a surname. comparative of cordy.
inion There is no word "inion" in the English language. It may be a misspelling of "onion," which is a type of vegetable. The inion is a term used in anatomy that refers to the most prominent projection of the occipital bone at the lower rear part of the skull in humans and other hominids. the external occipital protuberance of the skull.
Table 9: Representative examples of the words whose GPT definition embeddings are far from online dictionary definition embedding.
6 Word Embeddings and Definitions
Word embeddings are vector representations that capture the semantics of word usage. What dictionary definitions are for people, word embeddings are for NLP models: an easily-understood representation of the meaning of a vocabulary word. In this section we will directly compare traditional word embeddings to text embeddings of explicit dictionary embeddings, to help establish the level of correspondence between these representations.
Specifically, we obtain GloVE and 300-dimensional FastText word embeddings on all the words, and compare the closest neighbors of a word based on definition and word embeddings. We cannot compare these embeddings directly, because the underlying spaces and even dimensionality are incomparable.
Hence, we propose a new technique to measure the consistency of word and definition embeddings over a common vocabulary. For each word $w$ in the lexicon, we calculate the distance from $w$ to all other words, separately in word embedding and definition space. The correlation coefficent $r$ between these distances defines the agreement between the spaces from the perspective of word $w$. To get a full vocabulary metric on space similarity, we average these correlations for each word in the lexicon.
6.1 Frequency Analysis
Table 10 compares published and generated definitions with popular word embeddings, GloVE and FastText, using the correlation measure defined above. We distinguish between our three tranches of words, high-frequency, moderate, and rare.
The results are consistent across all sources and embeddings: high-frequency words show better consistency between definition-word embeddings than moderate-frequency words, which are substantially better than low frequency words. As our results in Section 5.2 show the definitions of roughly equal quality across frequency tranches, the difference must be due to the word embeddings themselves becoming less accurate with diminishing usage frequency. This presumably results from embeddings that are trained on insufficient amounts of data.
Model Frequency Glove FastText
GPT3-1 high 0.27 0.32
medium 0.21 0.28
low 0.08 0.22
GPT3-2 high 0.28 0.35
medium 0.22 0.30
low 0.13 0.24
GPT4-1 high 0.28 0.32
medium 0.23 0.30
low 0.11 0.25
GPT4-2 high 0.30 0.38
medium 0.26 0.35
low 0.12 0.29
Wordnet high 0.22 0.26
medium 0.19 0.22
low 0.11 0.16
Merriam-Webster high 0.25 0.30
medium 0.18 0.23
low 0.06 0.18
Dictionary.com high 0.33 0.31
medium 0.21 0.25
low 0.08 0.20
Table 10: Pearson correlation between word embedding and definition embedding based on cosine distance for different word frequency.
7 Conclusion and Future Work
The relationship between full text descriptions and embeddings is an interesting one. Our work demonstrates that model-generated definitions are accurate and lead to other useful representations.
Our experiments to detect long common phrases in alternate definitions of a particular word motivates questions of how long such phrases must be in generated text. How just how far apart can two texts be in surface form while still meaning the same thing? A study of situations where LLMs are highly constrained to specific word/phrase choices may send interesting light on the situation.
Limitations
The generative models powering ChatGPT are undergoing rapid evolution, so the observations in this paper cannot be guaranteed to hold in future releases. However, we do anticipate that the models will only continue to improve, and since our results show that they already exhibit similar accuracy and originality as published dictionaries we would be surprised of our observations do not continue to hold with future releases.
Having established basic agreement between publishing and generated definitions, as we have done in this paper, is a less demanding task than validating the completeness of multiple senses or alternate usages. We also have restricted our attention to English words and definitions. It would be interesting to evaluate model fidelity in other languages, or translation (e.g. “Define the French word rouge.”).
Appendix A Appendix
Model Definition of "wasp"
GPT3-1 a winged insect with a slender, often brightly colored body, the females of which have a poisonous sting
GPT3-2 A wasp is a predatory flying insect with a narrow waist and a sting, belonging to one of several families of the order Hymenoptera, such as the Vespidae, which includes the social wasps and hornets, or the Ichneumonidae, which includes parasitic wasps.
GPT4-1 "Wasp" has a couple of different meanings, depending on the context: 1. In entomology (the study of insects), a wasp is a type of flying insect, known for its black and yellow body, narrow waist, and ability to sting. There are many thousands of species of wasps, and they are found all over the world. 2. In social contexts, especially in the United States, WASP is an acronym for White Anglo-Saxon Protestant. It refers to a socio-economic and cultural group that has traditionally been dominant in American society, particularly among the upper class. This usage can sometimes carry implications about privilege and exclusivity. Remember that the context usually makes clear which meaning is intended.
GPT4-2 A wasp is a type of insect that is part of the order Hymenoptera and the suborder Apocrita, which is not a true ant or a bee. There are many different species of wasps, which can vary significantly in terms of size, color, and behavior. Most wasps are known for their ability to sting in self-defense. Unlike bees, which typically can sting only once before dying, many wasps can sting multiple times. Wasp stings are painful and can be dangerous to people who are allergic to them. Wasps can be categorized as either social or solitary. Social wasps, like the common yellowjacket, live in large colonies, and their nests can be found in trees, roof eaves, or underground. Solitary wasps, on the other hand, live alone and do not form colonies. Wasps play an important role in the ecosystem, as they prey on other pests, helping to control their populations. Some species are also important pollinators.
WordNet white Anglo-Saxon Protestant (a white person of Anglo-Saxon ancestry who belongs to a Protestant denomination); (social or solitary hymenopterans typically having a slender body with the abdomen attached by a narrow stalk and having a formidable sting)
Merriam-Webster any of numerous social or solitary winged hymenopterous insects…; any of various hymenopterous insects…; an American of Northern European and especially British ancestry and of Protestant background, especially : a member of the dominant and the most privileged class of people in the U.S.
Dictionary.com any of numerous social or solitary hymenopterous insects…; a person who is snappish or petulant; a member of the Women’s Air Force Service Pilots, an organization disbanded in 1944; a white Anglo-Saxon Protestant; a member of the privileged, established white upper middle class in the U.S.
Table 11: Definitions of the word “wasp” from GPT models and online dictionaries.